# Quantitative entropy estimates for 2D stochastic vortex model on the whole space under moderate interactions

### 05 Mar 2026 v3

@@ Abstract
We derive quantitative estimates for the stochastic 2D vortex model on the whole Euclidean space from stochastic particle systems driven by individual and environmental noises, obtaining pathwise quantitative bounds in terms of relative entropy. The main novelties lie in the application of the Donsker–Varadhan inequality in the framework of moderately interacting particle systems to handle the nonlinear terms, as well as in the introduction of localization techniques combined with the probabilistic data setting, in order to derive estimates for the quadratic variation terms. As an application of our entropy bounds, we derive a new energy estimate for the distance between the regularized empirical measure and the solution to the aforementioned model by combining the control of the Fisher information of the particle system, the Ladyzhenskaya inequality, and a nonlinear Grönwall lemma. Moreover, we prove the existence of a suitable solution to the limiting process.
@@ 1 Introduction
In this work, we derive quantitative entropy estimates for the 2D stochastic vortex model in the whole Euclidean space, which is a particular case of the following stochastic Fokker–Planck equation
$\mathrm{d}\rho_{t}=\Delta\rho_{t}\,dt+\frac{1}{2}D^{2}\rho_{t}(\sigma\sigma^{\top})_{t}\,dt-\nabla\cdot(\rho_{t}(K\ast\rho_{t}))\,dt-\nabla\rho_{t}\cdot\sigma_{t}\,dB_{t},$ (1)
@@
To achieve these bounds, we study the asymptotic behavior, as $N\to\infty$ of the empirical measure defined by
$S_{t}^{N}\doteq\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}},\qquad t\geqslant 0,$ (3)
where $\delta_{a}$ is the Dirac delta measure concentrated at $a\in\mathbb{R}^{d}$, by introducing the mollified empirical measure
@@
which allows us to apply the relative entropy functional directly to the densities $\rho^{N}$ and $\rho$.
To the best of our knowledge, this is the first time that pathwise quantitative bounds, in the sense of the relative entropy functional, have been derived for the 2D stochastic vortex model in the whole space via moderately interacting particles. This extends previously known results for bounded kernels on $\mathbb{R}^{d}$ [33] and for sub-Coulombian kernels under periodic boundary conditions [40]. As a consequence of our entropy bounds, we obtain energy estimates for the difference between the regularized empirical measure and the solution of (1), extending previous results by [27] and [37].
Our main novelty lies in the key technical tools we develop. For the first time in the context of moderately interacting particles, we apply the Donsker–Varadhan inequality, together with the dissipation provided by the Fisher information, to handle the nonlinearity. In addition, we employ localization techniques combined with our probabilistic data setting to derive estimates for the quadratic variation terms in the whole space, since the estimates established in [40] under periodic boundary conditions do not apply in the present framework. Finally, new energy estimates quantifying the distance between the regularized empirical measure and the solution to (1) are derived by combining the control of the Fisher information of the particle system with the Ladyzhenskaya inequality and a nonlinear Grönwall inequality. Moreover, we prove the existence of a suitable solution to the equation (1).
@@
Quantitative estimates for singular interactions have been substantially developed in recent years. The relative entropy method for establishing quantitative propagation of chaos in McKean–Vlasov systems with singular interaction kernels was first introduced in [24]. It was developed in the context of general first-order systems with interaction kernels in $W^{-1,\infty}$, a framework that includes the 2D viscous vortex model. We also refer to [43], where quantitative propagation of chaos for McKean–Vlasov equations with singular interaction kernels are established via the modulated energy method. In contrast to the relative entropy approach, which operates at the level of the joint law of the particle system, the modulated energy method is formulated directly in terms of the empirical measure associated with the particles.
Since then, the aforementioned works have paved the way for establishing quantitative propagation of chaos results. In [2] and [3], the strategies given in [24] and [43], were adapted to obtain a quantitative derivation for the Patlak–Keller–Segel model, in optimal subcritical regimes. Additionally, [21], by means of a logarithmic Sobolev inequality, derived uniform in time propagation of chaos for the 2D vortex model, adapting the strategy in [24]. Alternatively, with sharp estimates, global in time derivation of that model, based on relaxation estimates for the limiting equation and inequalities of Riesz modulated energy, was derived by [12], see also [11], [32] and [42], for the modulated energy method. Regarding sub-Coulombian kernels, [6] obtained the derivation of the mean-field approximation for Landau-like equations, while [41] obtained global-in-time mean-field convergence for gradient (conservative) diffusive flows. By extending the BBGKY hierarchy introduced in [31] to the case of singular kernels, and adapting the strategy developed in [24], [48] obtained the optimal convergence rate for the relative entropy between the k-particle marginals and the corresponding tensorized law.
@@
Quantitative estimates in the whole space were derived, in the context of moderately interacting particles  [^2] The moderately interacting particle model was introduced by Oelschläger in [35, 34, 36]. with singular kernels and without common noises - including the Biot-Savart, the Keller-Segel and the Dirac delta measure kernels - by [37] and [38], based on a semigroup approach introduced in [18], with further advancements in [8], [22], [28] and [39]. It is important to note that these works do not rely on the relative entropy method, but instead establish quantitative estimates in strong functional topologies. The connection between the time evolution of the relative entropy, of the joint distribution, and moderate interacting regime, on the whole space was carried out by [10], based on a combination between the relative entropy and the regularised $L^{2}$-estimate in [36], deriving a propagation of chaos result for the viscous porous medium equation.
However, it was in [40] that equation (1) was derived on the torus for singular kernels and common noise in the moderately interacting regime, by applying the relative entropy directly to the mollified empirical measure. Earlier, also in the same domain, the results established in [37] were extended to the setting with environmental noise in [27]. It is worth highlighting, that quantitative entropy estimates for systems with individual and common noise, have been obtained, as shown in [44] for incompressible the Navier-Stokes equations, [9] for the Hegselmann-Krause model, and [33] for mean-field systems with bounded kernels. For more results, see [19], [15], [16], and [29]. Particle systems driven by environmental noises, such as given by (2), appear in several contexts, such as problems in statistical mechanics, population dynamics and biology, as well as in the analysis of numerical Monte–Carlo schemes and related applications; see, e.g., [4], [5], [25], [20], and [30].
In contrast to the aforementioned works, our application of the relative entropy method provides pathwise estimates, deriving bounds directly at the trajectory level of the particle system. In those works, the estimates are obtained at the level of the joint law.
@@
In order to present our results, this work is organized as follows. In the next two subsections, 1.1 and 1.2, respectively, we introduce the notations and assumptions used throughout the text, setting the stage for the subsequent discussions. In the subsection 1.3, we present our main results. The sections 2 and 3 are dedicated to the proofs of our main results. To ease the presentation, the Appendix, provides the existence of a suitable solution for the limiting equation, as well as some inequalities used throughout the proofs.
@@ 1. $(\mathbf{A}^{V})$
    Let $V^{N}:\mathbb{R}^{d}\to\mathbb{R}$ be the scaling given by $V^{N}(y)=N^{\beta}V(N^{\frac{\beta}{d}}y)$, $\beta\in(0,1)$, with $V\in\mathcal{P}\left(\mathbb{R}^{d}\right)$ satisfying
@@
    $V(y)\leq C_{d}(1+|y|^{2})^{-\alpha},$ (6)
    and for $\gamma\in(0,1)$ and $y\neq 0$
    $V(y)\leq C_{d}|y|^{-3d-\gamma}.$ (7)
@@ 2. $(\mathbf{A}^{K})$
    It holds for $\gamma\in(1/2,1)$ that
    $\|K\ast f\|_{\gamma}\leq C\left(\|f\|_{1}+\|f\|_{4}\right)$ (8)
    for all $f\in H^{1}$.
@@ 5. $(\mathbf{A}^{\rho_{0}})$
    The initial condition $\rho_{0}$ is taken such that $\rho_{0}\in W^{2,\infty}\cap\mathcal{P}\left(\mathbb{R}^{d}\right)$ and
@@ 6. $(\mathbf{A}^{X_{0}})$
It follows that, this singular kernel satisfies the Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$. Furthermore, $K$ satisfy $(\mathbf{A}^{K})$, see [39].
@@
Assumption $(\mathbf{A}^{\rho_{0}})$ was formulated in [7, 49, 50], where, relying on Li–Yau–type gradient bounds and Hamilton–type heat kernel estimates, the authors use it to control the solution and its derivatives up to second order. It allows the relative entropy approach of [24] to be extended from the periodic framework to the whole space.
1.3 Statement of the main results
In this subsection, we establish our main results. We begin by formulating a suitable notion of solution, needed for deriving quantitative entropy estimates.
@@ 1.3 Statement of the main results
Assume $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, and $(\mathbf{A}^{\rho_{0}})$. A pathwise solution for (1) with initial condition $\rho_{0}$, is a measurable function $\rho:[0,T_{1}]\times\mathbb{R}^{d}\times\Omega\to\mathbb{R}$, for some $T_{1}>0$, such that:
@@
    $\rho\in L^{\infty}\left([0,T_{1}];W^{2,1}\cap W^{2,\infty}\cap\mathcal{P}(\mathbb{R}^{d})\right)$ (15)
@@
    for all $t\in[0,T_{1}]$ and constants $C_{1},C_{2}$ and $C_{3}$ dependent on $\widetilde{C}_{1}$, $\widetilde{C}_{2}$ and $\widetilde{C}_{3}$.
We are now able to extend what was known before for bounded kernels in the whole space as in [33] and in [40], for sub-Coulombian kernels, under periodic boundary conditions. Our main result can be stated as follows.
@@
$\theta=\min\left(1-\beta(1+\frac{2}{d}+2\alpha);\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta$
@@
Then, there exist a time $T\in(0,T_{1})$, an almost surely finite random variable $A_{0}$, and a set $\Lambda\in\mathcal{F}$, $\mathbb{P}(\Lambda)=1$, with the following property: for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (21)
@@
The following result establishes energy estimates as an application of our entropy bounds, extending previous results by [27] and [37]. Precisely, we utilize the entropy dissipation to control the Fisher information of the particle system, which in turn allows us to derive, a new energy estimate for the difference between the regularized empirical measure and the solution to (1), using the Ladyzhenskaya inequality and a nonlinear Gronwall lemma.
Theorem 2.
$\lim_{N\to\infty}N^{\widetilde{\theta}}\mathcal{H}(\rho_{0}^{N}|\rho_{0})=0,\,\,\,\mathbb{P}-a.s.,$ (23)
$\lim_{N\to\infty}N^{\widetilde{\theta}}\left\|\rho_{0}^{N}-\rho_{0}\right\|_{2}^{2}=0,\,\,\,\mathbb{P}-a.s.$ (24)
and
$\widetilde{\theta}\doteq\min{\left(\theta_{1},\theta_{2},\widetilde{\theta}_{1},\widetilde{\theta}_{2},\widetilde{\theta}_{3},\widetilde{\theta}_{4}\right)}-\delta$
where
$\theta_{1}\doteq\left(1-\beta\left(2+2\alpha\right)\right);\,\,\,\theta_{2}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{2}\Big)\right);\,\,\,\widetilde{\theta}_{1}\doteq\left(1-2\beta(1+\frac{1}{2})+\beta\right)$
$\widetilde{\theta}_{2}\doteq\beta(2-3\gamma);\,\,\,\widetilde{\theta}_{3}\doteq\frac{\beta\gamma^{2}}{2};\,\,\,\widetilde{\theta}_{4}\doteq\left(\frac{1}{2}-\beta\Big(1+\frac{1}{2}\Big)\right)$
with $\gamma\in(1/2,2/3)$ and $\delta>0$, such that $\widetilde{\theta}>0$ and $\beta\in\left(0,\frac{1}{2+2\alpha}\right)$.
Then, there exist a time $T\in(0,T_{1})$, a pair of almost surely finite random variables $A_{0}$, $\widetilde{A}_{0}$, and a set $\Lambda\in\mathcal{F}$, $\mathbb{P}(\Lambda)=1$, with the following property: for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have
$\sup_{t\in[0,T]}\|\rho_{t}^{N}-\rho_{t}\|_{2}^{2}+\int_{0}^{T}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds$
[…]
In particular, we get
$\limsup_{N\to\infty}\left(\sup_{t\in[0,T]}\|\rho_{t}^{N}-\rho_{t}\|_{2}^{2}+\int_{0}^{T}\left\|\nabla(\rho_{s}-\rho^{N}_{s})\right\|_{2}^{2}\,ds\right)$ $\lesssim T,\,\,\,\mathbb{P}-a.s.$ (26)
As a corollary of Theorem 1, we establish quantitative bounds for the marginals of the empirical measure of the particle system. In the sense of [1, Section 8.3], it corresponds to get bounds of the empirical measure in the sense to the Kantorovich–Rubinstein metric, given in (4).
Corollary 1.
Under the assumptions of Theorem 1, for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$, such that for $N\geq N_{0}$ we have
$\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim\left(\mathcal{H}(\rho_{0}^{N}|\rho_{0})+N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (27)
In particular, we get
$\limsup_{N\to\infty}\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $\lesssim T,\,\,\,\mathbb{P}-a.s.$ (28)
Remark 4.
We observe that, compared with previous works on moderate particle systems without common noise, we lose the order of convergence. This is because the estimation of the martingale term in our work is is of order $\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)$, while in [37] is the order $\frac{1}{2}-\frac{\beta}{2}$, see Theorem 1.3 in [37]. When environmental noise is taken into account, the overall framework remains comparable to that of [27]. For the same reason, the range of the parameter $\beta$ is smaller in our work. However, in the Theorem 1, we obtain an improved convergence rate compared to that established in [40].
@@
2 Proofs of Theorem 1
In this section we present the proofs of Theorem 1. In Subsection 2.1, by applying the Itô’s formula, such as in [40], we derive an evolution equation for the relative entropy functional of the regularized empirical measure with respect to the solution of (1). In Subsection 2.2, we deal with the quadratic variation terms arising from the application of Itô’s formula. In Subsection 2.3, we address the nonlinear contributions appearing in the evolution equation for the relative entropy functional obtained in the first step.
Finally, in Subsections 2.4 and 2.5, combining the previous estimates with Gronwall’s lemma and suitable localization techniques, we close the argument and complete the proof.

### 13 Apr 2026 v4

# Quantitative propagation of chaos for 2D stochastic vortex model on the whole space under moderate interactions

@@ Abstract
We derive the stochastic 2D vortex model on the whole Euclidean space from moderately interacting particle systems driven by individual and environmental noises, obtaining quantitative estimates in the sense of the entropy and energy functionals. The main novelties lie in combining the control of the Fisher information of the particle system with the Ladyzhenskaya and Donsker–Varadhan inequalities, as well as localization techniques within the probabilistic data setting, to address the nonlinearity and quadratic variation arising from Itô’s formula. Moreover, we construct a suitable solution for the limiting process.
@@ 1 Introduction
In this work, we establish a quantitative derivation of the stochastic 2D vortex model in the whole Euclidean space, which is a particular case of the following stochastic Fokker–Planck equation
$\mathrm{d}\rho_{t}=\Delta\rho_{t}\,dt+\frac{1}{2}\nabla^{2}\rho_{t}(\sigma\sigma^{\top})_{t}\,dt-\nabla\cdot(\rho_{t}(K\ast\rho_{t}))\,dt-\nabla\rho_{t}\cdot\sigma_{t}\,dB_{t},$ (1)
@@
To achieve it, we study the asymptotic behavior, as $N\to\infty$ of the empirical measure defined by
$S_{t}^{N}\doteq\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}},\qquad t\geqslant 0$
by introducing the mollified empirical measure
@@
which allows us to apply the entropy and energy functionals directly to the densities $\rho^{N}$ and $\rho$.
Particle systems driven by environmental noises appear in several contexts, such as in statistical mechanics, population dynamics, as well as in numerical analysis, see, e.g., [4], [5], [23], [18], and [27], and quantitative bounds characterize a reduction in the complexity of the system (2), for large values of $N$, see [22].
To the best of our knowledge, this is the first time that, propagation of chaos, supplemented with pathwise quantitative bounds, in the sense of the relative entropy functional, have been derived for the 2D stochastic vortex model in the whole Euclidean space via moderately interacting particles. This extends previously known results for bounded kernels on $\mathbb{R}^{d}$ [29] and for sub-Coulombian kernels under periodic boundary conditions [36]. In addition, we obtain energy estimates for the difference between the regularized empirical measure and the solution of (1), extending previous results by [25] and [33].
Our main novelty lies in the key technical tools we combine. Using Itô formula, we compute the evolution equations for the entropy and energy functionals associated with the regularized empirical measure and the solution of the aforementioned model. By combining the control of the Fisher information of the particle system with the Ladyzhenskaya and Donsker–Varadhan inequalities, along with localization techniques within the probabilistic data setting, to deal with the nonlinearity and quadratic variation, we are able to close the argument with a nonlinear Grönwall inequality. Moreover, we prove the existence of a suitable solution to equation (1).
@@
Quantitative estimates for singular interactions have been substantially developed in recent years. The relative entropy method for establishing quantitative propagation of chaos with singular interaction kernels in $W^{-1,\infty}$ was first introduced in [22], providing a framework that includes the 2D viscous vortex model. Alternatively, we refer to [39], which establishes quantitative propagation of chaos for McKean–Vlasov equations with singular interaction kernels via the modulated energy method. In contrast to the relative entropy approach, which operates at the level of the joint law of the particle system, the modulated energy method is formulated directly in terms of the empirical measure associated with the particles.
Since then, the aforementioned works have paved the way for establishing quantitative propagation of chaos results, under singular interactions. In [2] and [3], the strategies given in [22] and [39], were adapted to obtain a quantitative derivation for the Patlak–Keller–Segel model, in optimal subcritical regimes. Additionally, [19], by means of a logarithmic Sobolev inequality, derived uniform in time propagation of chaos for the 2D vortex model, adapting the strategy in [22]. Alternatively, with sharp estimates, global in time derivation of that model, based on relaxation estimates for the limiting equation and inequalities of Riesz modulated energy, was derived by [12], see also [11], [28] and [38], for the modulated energy method. Regarding sub-Coulombian kernels, [6] obtained the derivation of the mean-field approximation for Landau-like equations, while [37] obtained global-in-time mean-field convergence for gradient (conservative) diffusive flows. Moreover, quantitative entropy estimates for systems with individual and common noise, have been obtained, as shown in [40] for the 2D vortex model, on torus, [9] for the Hegselmann-Krause model.
@@
Quantitative estimates in the whole space were derived by [33] and [34] in the context of moderately interacting particles [^2] The moderately interacting particle model was introduced by Oelschläger in [31, 30, 32]. without common noises and with singular kernels (including the Biot-Savart, Keller-Segel, and Dirac delta measure kernels). Their derivation is based on a semigroup approach introduced in [17] and has motivated further advancements, for example, in [8, 15, 20, 26, 35]. Under periodic boundary conditions, [25] used the Krilov $L^{p}$ theory for SPDE’s, to extend the results given in [33], in the presence of environmental noises, as the semigroup approach is not available in this setting. It is important to note that these works do not rely on the relative entropy method, but instead establish quantitative estimates in strong functional topologies.
The connection between the time evolution of the relative entropy, of the joint distribution, and moderate interacting regime, on the whole space was carried out by [10], based on a combination between the relative entropy and the regularised $L^{2}$-estimate in [32], deriving a propagation of chaos result for the viscous porous medium equation. However, it was in [36] that equation (1) was derived on the torus for singular kernels and common noise in the moderately interacting regime, by applying the relative entropy directly to the mollified empirical measure.
In this work, we extend the quantitative entropy estimates given in [36] under periodic boundary conditions and in [29] for bounded kernels to the whole space for the stochastic 2D vortex model. Additionaly, we obtain energy estimates for that model, extending the results in [25] and [33]. In contrast to the aforementioned works, which establish quantitative entropy estimates at the level of the joint law, our application of the relative entropy method provides pathwise estimates, yielding bounds directly at the trajectory level of the particle system.
@@
In order to present our results, this work is organized as follows. In the next three subsections, 1.1, 1.2, and 1.3, we introduce the notations and assumptions used throughout the text, setting the stage for the subsequent discussions, as well as the notion of solution to (1). In the subsection 1.4, we present our main results. The sections 2 and 3 are dedicated to the proofs of our main results. Finally, the Appendix provides the existence of a suitable solution for the limiting equation, as well as some inequalities used throughout the proofs.
@@ 1. $(\mathbf{A}^{V})$
    Let $V^{N}:\mathbb{R}^{d}\to\mathbb{R}$ be the scaling given by $V^{N}(y)\doteq N^{\beta}V(N^{\frac{\beta}{d}}y)$, $\beta\in(0,1)$, with $V\in\mathcal{P}\left(\mathbb{R}^{d}\right)$ satisfying
@@
    $V(y)\leq C_{d,\alpha}(1+|y|^{2})^{-\alpha},$ (5)
    and for $\gamma,\beta\in(0,1)$ and $y\neq 0$
    $V(y)\leq C_{d,\gamma,\beta}|y|^{-3d/\beta-\gamma}.$ (6)
@@ 2. $(\mathbf{A}^{K})$
    It holds for $\gamma\in(0,1/2)$ that
    $\|K\ast f\|_{\gamma}\leq C_{\gamma}\left(\|f\|_{1}+\|f\|_{4}\right)$ (7)
    for all $f\in W^{1,2}$.
@@ 5. $(\mathbf{A}^{\rho_{0}})$
    The initial condition $\rho_{0}$ is taken such that $\rho_{0}\in W^{2,1}\cap W^{2,\infty}\cap\mathcal{P}\left(\mathbb{R}^{d}\right)$, and
@@ 6. $(\mathbf{A}^{X_{0}})$
It follows that, this singular kernel satisfies the Assumption $\left(\mathbf{A}^{\nabla\cdot K}\right)$. Furthermore, $K$ satisfies $(\mathbf{A}^{K})$, see [35], page 8.
@@
Assumption $(\mathbf{A}^{\rho_{0}})$ was formulated in [7, 44, 45], where, relying on Li–Yau–type gradient bounds and Hamilton–type heat kernel estimates, the authors use it to control the solution and its derivatives up to second order. It allows the relative entropy approach of [22] to be extended from the periodic framework to the whole space. It motivates the following definition.
1.3 Solution of limiting process
We now formulate a suitable notion of solution of (1) to derive our quantitative propagation of chaos.
@@ 1.3 Solution of limiting process
Assume $(\mathbf{A}^{\nabla\cdot K})$, $(\mathbf{A}^{\sigma})$, and $(\mathbf{A}^{\rho_{0}})$. A pathwise solution for (1) with initial condition $\rho_{0}$ is a measurable function $\rho:[0,T]\times\mathbb{R}^{d}\times\Omega\to\mathbb{R}$, for some $T>0$, such that:
@@
    $\rho\in L^{\infty}\left([0,T];W^{2,1}\cap W^{2,\infty}\cap\mathcal{P}(\mathbb{R}^{d})\right)$ (14)
@@
    for all $t\in[0,T]$ and constants $C_{1},C_{2}$ and $C_{3}$ dependent on $\widetilde{C}_{1}$, $\widetilde{C}_{2}$ and $\widetilde{C}_{3}$.
1.4 Statement of the main results
We are now able to extend what was known before for bounded kernels in the whole space as in [29] and in [36], for sub-Coulombian kernels, under periodic boundary conditions, by deriving entropy estimates. Also we extend previous results by [25] and [33], by establishing energy estimates. Our main result is stated as follows.
Theorem 1.
$\lim_{N\to\infty}N^{\theta}\mathcal{H}(\rho_{0}^{N}|\rho_{0})=0,\,\,\,\mathbb{P}-a.s.,$ (19)
$\lim_{N\to\infty}N^{\theta}\left\|\rho_{0}^{N}-\rho_{0}\right\|_{2}^{2}=0,\,\,\,\mathbb{P}-a.s.$ (20)
and
$\theta\doteq\min{\left(\left(1-\beta\left(2+2\alpha\right)\right);\left(\frac{1}{2}-\beta\Big(1+\frac{1}{2}\Big)\right);\frac{\beta\gamma^{2}}{2}\right)}-\delta$
with $\gamma\in(0,1/2)$ and $\delta>0$, such that $\theta>0$ and $\beta\in\left(0,\frac{1}{4}\right)\cap\left(0,\frac{1}{\big[2+2\alpha\big]}\right)$.
Then, there exists a time $T>0$, such that
$\lim_{N\to\infty}N^{\theta}\left(\sup_{t\in[0,T]}\left(\mathcal{H}(\rho_{t}^{N}|\rho_{t})+\left\|\rho_{t}^{N}-\rho_{t}\right\|_{2}^{2}\right)+\int_{0}^{T}\left\|\nabla(\rho_{s}^{N}-\rho_{s})\right\|_{2}^{2}\,ds\right)$ $=0,$ (21)
almost surely.
As a corollary of Theorem 1, we establish quantitative propagation of chaos for the marginals of the empirical measure of the particle system. In the sense of [1, Section 8.3], it corresponds to get convergence of the empirical measure in the sense to the Kantorovich–Rubinstein metric, given in (3).
Corollary 1.
Under the assumptions of Theorem 1, we have
$\lim_{N\to\infty}N^{\theta}\sup_{t\in[0,T]}\left\|S_{t}^{N}-\rho_{t}\right\|_{0}^{2}$ $=0,\,\,\,\mathbb{P}-a.s.$ (22)
In dimensions greater than or equal to two, we establish the boundedness of the relative entropy functional:
@@ 1.4 Statement of the main results
$\theta=\min\left(1-\beta\left(1+\frac{2}{d}+2\alpha\right);\frac{1}{2}-\beta\Big(1+\frac{1}{d}\Big)\right)-\delta$
@@
Then, there exist a time $T>0$, an almost surely finite random variable $A_{0}$, and a set $\Lambda\in\mathcal{F}$, $\mathbb{P}(\Lambda)=1$, with the following property: for all $\omega\in\Lambda$, there exists $N_{0}=N_{0}(\omega)$ such that, if $N\geq N_{0}$ we have
$\sup_{t\in[0,T]}\mathcal{H}(\rho_{t}^{N}|\rho_{t})$ $\lesssim\left(N^{-\theta}+A_{0}N^{-\theta}+T\right).$ (24)
@@
2 Proof of Theorem 1
In this section we present the proofs of Theorem 1. In Subsection 2.1, by applying the Itô’s formula, such as in [36], we derive an evolution equation for the relative entropy functional of the regularized empirical measure with respect to the solution of (1). In Subsection 2.1.1, we deal with the quadratic variation terms arising from the application of Itô’s formula, while in Subsection 2.1.2, we address the nonlinear contributions appearing in the evolution equation for the relative entropy functional obtained in the first step.
In Subsection 2.2, we derive an evolution equation for the energy, relating our densities, being the nonlinear term, arising from this computation, estimated in Subsection 2.2.1, as in Subsection 2.3, we combine our entropy and energy estimates.
Finally, in Subsection 2.4, by a nonlinear Gronwall Lemma and suitable localization techniques, we close the argument and complete the proof.